OpenStudy (anonymous):

TW is a diameter of circle X, and TW is parallel to UV...... If the measure of is 25°, what is the degree measure of ?

OpenStudy (anonymous):



OpenStudy (raden):

what is the degree measure of "what" ?

OpenStudy (anonymous):

It's just asking what the degree measures are. The answer choices are ( 210, 130, 155, 115)

OpenStudy (anonymous):

What the degree measure is*

OpenStudy (raden):

25 is the measure arc TU, is your question asking the measure of arc UW ?

OpenStudy (anonymous):

My bad, I didn't realize I messed that question up. It is supposed to say " If the measure of is TU 25°, what is the degree measure of UV ?

OpenStudy (anonymous):

TW*

OpenStudy (raden):

the measure of UV is  UV = 180 - (25+25) = 180 - 50 = 130